Rowena wants to test a randomly selected 20 students for her study. Identify the t value of the following percentiles.



1. 75th percentile


2.90th percentile


3. 60th percentile.


4. 99th percentile


5.50th percentile

1
Expert's answer
2022-04-12T17:46:07-0400

We use the t-table, degrees of freedom df = n - 1 = 20 - 1 = 19.

  1. 75th percentile = 0.687621
  2. 90th percentile = 1.327728
  3. 60th percentile = 0.256923
  4. 99th percentile = 2.239483
  5. 50th percentile = 0